OpenStudy (ttp):

Which grouping of numbers gives the expression a value of 315? the product three times ten divided by five minus three plus three hundred three times the quotient ten divided by five minus three plus three hundred three times ten divided by the difference five minus three plus three hundred three times ten divided by five minus the quantity three plus three hundred

OpenStudy (texaschic101):

1.) (3 x 10) / 5 - 3 + 300 30/5 - 3 + 300 6 - 3 + 300 3 + 300 = 303 (not it) 2.) 3 * (10/5 ) - 3 + 300 3 * 2 - 3 + 300 6 - 3 + 300 3 + 300 = 303 (nope) 3.) (3 * 10) / (5 - 3) + 300 30/2 + 300 15 + 300 = 315 (BINGO) 4.) (3* 10)/5 - (3 + 300) 30/5 - 303 6 - 303 = - 297 (nope)

OpenStudy (ttp):

Thanks:) again:)